OpenStudy (anonymous):

Fill in the missing number. ? x 30 = 10 ? x 100 = 50 ? x 100 = 75

OpenStudy (anonymous):

They are all seperate by the way. Not one whole problem. :)

OpenStudy (mertsj):

What is the inverse of multiplication?

OpenStudy (anonymous):

I don't know?

OpenStudy (mertsj):

How do you "undo" multiplication?

OpenStudy (unklerhaukus):

division

OpenStudy (mertsj):

Unkle wants to help you now.

OpenStudy (anonymous):

Someone please help me!

OpenStudy (unklerhaukus):

\[ ? \times 30 = 10\]\[? = 10/30\]\[?=1/3\]

OpenStudy (unklerhaukus):

? x 100 = 50 ? = 50/100 ?=5/10 ?=1/2

OpenStudy (unklerhaukus):

try then next one ? x 100 = 75 divide both sides by 100

OpenStudy (anonymous):

1/75?

OpenStudy (unklerhaukus):

\[? \times 100 \div 100= 75 \div 100\]

OpenStudy (unklerhaukus):

\[?={75\over 100}\] \[?={3 \over 4}\]

OpenStudy (anonymous):

Than answers are in the attachment.